OpenStudy (anonymous):

du(x)/dx- tan(x)u(x)=-2sinx find u(x) inhomogeneous differential equation

OpenStudy (melissa_something):

lol wtf

OpenStudy (anonymous):

actually im trying to solve inhomogeneous differential equation

OpenStudy (anonymous):

du(x)/dx- tan(x)u(x)=-2sinx

OpenStudy (melissa_something):

Is that calculus or..?

OpenStudy (anonymous):

advance applied mathematics

OpenStudy (anonymous):

I think the integral you looking for is: \[e^{\int\limits\tan x}\] which would be integrating factor for this equation

OpenStudy (anonymous):

well the answer appears to be u(x)=c(1/cos(x))+cos(x)

OpenStudy (anonymous):

but i dont know how to get to there

OpenStudy (anonymous):

I am a bit rusty on this, but here it goes (as much as I remmeber) 1º Find the integrating factor of the 1º order linear nonhomogeneous differential equation u'-Pu=Q, which is e^{int P}. Inthis case it is equal 1/cos x. 2º Multiply bouth side of the equation by this factor. u'/cosx -u(tanx/cosx)=-2sinx/cosx 3º Notice that the left side is derivative of u/cosx, so d(u/cosx)/dx = -2tanx 4º Integrate bouth sides: u/cosx=-2 int tanx

OpenStudy (anonymous):

I think I made a mistake in integrating factor is not 1/cosx, but just cosx and the answer should be u=2+c/cosx I think

OpenStudy (anonymous):

just tried but it dint give the desired answer...

OpenStudy (anonymous):

u(x)=c(1/cos(x))+cos(x) this was the given answer

OpenStudy (anonymous):

Integrating factor = cosx after multiply: cosx u' -sinx u =-2sinx cosx d(cosx u)/dx = - 2sinx cosx cosx u =-2 int {sinx cosx dx } cosx u = 2 cos x + C u=2 +C/cosx

OpenStudy (anonymous):

yeah im getting the same answer as yours..im wondering why does the book is written as u=cosx +C/cosx

OpenStudy (anonymous):

maybe I made a mistake somewhere............

OpenStudy (anonymous):

hmm not so sure too...but thanks alot @myko